OpenStudy (anonymous):

solve sq. y-13 + sq.y =13

myininaya (myininaya):

First get 1 sqrt thingy by itself then square both sides so you can get rid of that sqrt but you will still have another sqrt on the other side so you get that one by itself and then square again

myininaya (myininaya):

no more sqrt thingy

myininaya (myininaya):

so we have (y-13)^(1/2)=13-y^1/2

myininaya (myininaya):

square both sides

myininaya (myininaya):

y-13=(13-y^(1/2))^2

OpenStudy (anonymous):

sorry to interrupt. @myinaya, like the way you said "sqrt thingy" XD

myininaya (myininaya):

lol

myininaya (myininaya):

i just thought it is easier to understand

myininaya (myininaya):

y-13=169-2y^1/2+y

myininaya (myininaya):

remember get that last sqrt thingy by itself

myininaya (myininaya):

y-y-13-169=2y^(1/2)

myininaya (myininaya):

oh mine thought we ran into some like terms. Lets just combine them real quick

myininaya (myininaya):

though*

myininaya (myininaya):

-199=2y^(1/2)

myininaya (myininaya):

so we have -199/2=y^(1/2) then square both sides (-199/2)^2=y

myininaya (myininaya):

You should always check your solutions when you raise both sides to an even power though

OpenStudy (anonymous):

so y= -199/2^2

OpenStudy (anonymous):

y = (-199/2)^2

myininaya (myininaya):

and actually it was suppose to be (199/2)^2 but it doesnt matter since it is squared

OpenStudy (anonymous):

this is so confusing to me ....haven't been in school for 20+ years

myininaya (myininaya):

like i forgot to write a negative sign before the last step